# Towards Unified and Lossless Latent Space for 3D Molecular Latent Diffusion Modeling

### v3

@@ B.2 More Visualizations of Generated Molecules
Appendix C More Related Works
@@
Appendix D More Experimental Settings
D.1 Experimental Settings for De Novo 3D Molecule Generation
@@
Hyperparameters. 11 shows the key hyperparameters used for training the UAE-3D and UDM-3D models.
@@
Comparison to Other Evaluation Protocols. Discrepancies in evaluation methodologies can significantly impact reported performance metrics. For instance, ADiT’s validity scores stem from their reliance on pymatgen.core.Molecule for molecular reconstruction which prioritizes topological correctness over chemical stability. While this method validates basic atomic connectivity, it fails to rigorously assess stability and geometric feasibility as evidenced by ADiT’s poor performance on these metrics (Table 2). In contrast, our evaluation employs RDKit’s Chem.Mol for molecule reconstruction, which enforces stricter valency checks and chemical correctness. And our broad evaluation criteria comprehensively measure the validity, diversity, and geometric fidelity of 3D molecule generation.
D.2 Experimental Setting for Conditional 3D Molecule Generation
@@
1:  $\mathbf{Z}^{(1)}\sim\mathcal{N}(0,\mathbf{I})$Initial noise}
2:  for $k=K$ downto $1$ do
@@
12:  $\hat{\mathbf{G}}=\langle\hat{\mathbf{F}},\hat{\mathbf{E}},\hat{\mathbf{X}}\rangle\leftarrow\mathcal{D}(\mathbf{Z}^{(0)})$UAE-3D decoder}
D.3 Pseudo Code

### v4

@@ A.2 More Visualizations of Generated Molecules
Appendix B More Related Works
@@
Appendix C More Experimental Settings
C.1 Experimental Settings for De Novo 3D Molecule Generation
@@
Hyperparameters. 13 shows the key hyperparameters used for training the UAE-3D and UDM-3D models.
@@
Comparison to Other Evaluation Protocols. Discrepancies in evaluation methodologies can significantly impact reported performance metrics. For instance, ADiT’s validity scores are obtained by reconstructing molecules from atomic coordinates using pymatgen.core.Molecule, followed by bond inference via RDKit (likely using the xyz2mol heuristic [70]). This pipeline ensures topological consistency (i.e., SMILES validity) but does not explicitly enforce chemical stability constraints such as correct valency or bond geometry. In contrast, our evaluation employs RDKit’s Chem.Mol for molecule reconstruction, which enforces stricter valency checks and chemical correctness. And our broad evaluation criteria comprehensively measure the validity, diversity, and geometric fidelity of 3D molecule generation.
C.2 Experimental Setting for Conditional 3D Molecule Generation
@@
[…]
2:for $k=K$ downto 11 do
@@
[…]
C.3 Pseudo Code